# A preconditioned difference of convex functions algorithm with extrapolation and line search

### v1

@@ 4 Convergence analysis
This section analyzes the convergence of the two proposed algorithms and addresses a generalized special case. We first present the convergence for the case that $f(x)$ and $g_{1}(x)$ have Lipschitz gradients while $g_{2}$ is non-smooth. For the case $g_{1}$ being non-smooth, we rigorously prove convergence for the scenario where $g_{1}(x)$ is non-differentiable and $g_{2}(x)$ is Lipschitz differentiable under mild condition. This proof imposes standard assumptions on and selects a representative special case to demonstrate the core methodology. Collectively, these two cases can cover the most practical computational scenarios encountered in applications.
@@ 4.1 Convergence analysis of Algorithm 1
Let $\bar{x}^{n}$ is generated by (3.8), then the following inequality holds
@@
As $\bar{x}^{n}$ is a minimizer of the strongly convex function in (3.8) and by evaluating the function values at $\bar{x}^{n}$ and $x^{n}$, we obtain
$-\xi^{n},\bar{x}^{n}\right>+\frac{1}{2}\left\|\bar{x}-y^{n}\right\|_{M}^{2}+f\left(\bar{x}^{n}\right)+g_{1}\left(\bar{x}^{n}\right)\leq\left<-\xi^{n},x^{n}\right>+\frac{1}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
@@
$+f\left(x^{n}\right)+g\left(x^{n}\right)-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
@@
Remembering $y^{n}=x^{n}+\beta_{n}(x^{n}-x^{n-1})$, this leads to the inequality (4.1). ∎
@@
This leads to (4.3) and the proof is finished. ∎
@@
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimal condition.
@@
(iii). Given the boundedness of $\left\{x^{n}\right\}_{n}$ from (ii), there exists at least one cluster point $x^{*}$ with a convergent subsequence $\left\{x^{n_{i}}\right\}$ satisfying
@@
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)-\nabla f\left(\bar{x}^{n_{i}}\right)-\nabla g_{1}\left(\bar{x}^{n_{i}}\right)+\nabla(f+g_{1})({x}^{n_{i}})\in-\partial g_{2}({x}^{n_{i}})+\nabla(f+g_{1})({x}^{n_{i}}).$
@@
$-M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+(\nabla f({x}^{n_{i}})-\nabla f(\bar{x}^{n_{i}}))+(\nabla g_{1}({x}^{n_{i}})-\nabla g_{1}(\bar{x}^{n_{i}})).$ (4.13)
@@
Setting $\tilde{L}=L+L_{g_{1}}$, with $\bar{x}^{n}-x^{n}=\frac{\lambda_{n}}{1+\lambda_{n}}\left(x^{n+1}-x^{n}\right)$, $\bar{x}^{n-1}-x^{n-1}=\frac{\lambda_{n-1}}{1+\lambda_{n-1}}\left(x^{n}-x^{n-1}\right)$, we obtain the following formulation
@@
$\geq\psi\prime\left(A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)\cdot\left[A\left(x^{n+1},\bar{x}^{n},x^{n}\right)-A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)\right]\cdot$
@@
Finally, we obtain $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$. ∎
@@ 4.2 Convergence analysis of Algorithm 2
In contrast to Algorithm 1, Algorithm 2 requires the function $f$ to be convex. The following Lemma 3, Lemma 4 and Theorem 2 are obtained on this basis, and we will not write out one by one later on. Following the analytical framework established for Algorithm 2, we initially demonstrate that Lemma 3 serves to rigorously characterize the quasi-descent property intrinsic to the energy functional governing the system.
@@
Let $\bar{x}^{n}$ is generated by (3.10), then the following inequality holds
@@
$\nabla f\left(y^{n}\right)-\xi^{n},y\right>+\frac{L}{2}\left\|y-y^{n}\right\|_{M}^{2}+g_{1}\left(y\right).$
@@
$\nabla f\left(y^{n}\right)-\xi^{n},\bar{x}^{n}\right>+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(\bar{x}^{n}\right)\leq\left<\nabla f\left(y^{n}\right)-\xi^{n},x^{n}\right>+\frac{L}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
@@
$\nabla f\left(y^{n}\right),\bar{x}^{n}-x^{n}\right>+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(\bar{x}^{n}\right)-g_{2}\left(\bar{x}^{n}\right)\leq\frac{L}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
@@
Applying the matrix dominance $M\succeq I$ and $y^{n}=x^{n}+\beta_{n}(x^{n}-x^{n-1})$, we derive (4.18). ∎
@@
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is the critical point of the problem (1.1), i.e. it satisfies the first order optimal condition.
@@ 4.3 Convergence analysis for the special case with nonsmooth $g_{1}$
for any $\bar{x}\in\operatorname{dom}g_{1}$ and $\tilde{w}\in\partial g_{1}(\bar{x})$ and $\mu$ can be zero while $g_{1}$ is not strongly convex. We propose a more general case under the following Assumption 1. It can cover the widely used case $g_{1}=\|x\|_{1}$ along with $\nabla g_{2}(x)$ being Lipschitz continuous with constant $L_{g_{2}}$, which includes some important regularization such as SCAD as discussed in [1].
Assumption 1.
We assume the proper, lower semicontinuous and convex function $g_{1}$ has the following property, for any $\bar{x}\in\operatorname{dom}g_{1}$, any sufficiently small $\varepsilon>0$, and $x\in B_{\varepsilon}(\bar{x})\cap\operatorname{dom}g_{1}$, there exits a subgradient $\bar{w}\in\partial g_{1}(\bar{x})$ and $\sigma>0$ such that
$g_{1}\left(x\right)\leq g_{1}\left(\bar{x}\right)+\langle x-\bar{x},\bar{w}\rangle+\frac{\sigma}{2}\left\|x-\bar{x}\right\|^{2}.$ (4.21)
In the following proposition, we rigorously demonstrate that the $l_{1}$-norm satisfies Assumption 1. This analysis confirms the theoretical soundness of our proposed algorithmic framework and verifies that the required assumption is readily satisfiable in practical implementations.
Proposition 3.
The anisotropic $l_{1}$-norm in $\mathbb{R}^{n}$
$h\left(x\right)=\left\|x\right\|_{1}=\left|x_{1}\right|+\left|x_{2}\right|+\ldots+\left|x_{n}\right|$
satisfies the Assumption 1.
Proof.
Let us first consider the case $h(x)=|x|$ with $x\in\mathbb{R}$. It is known that
$\partial h(x)$ =$=\begin{cases}1&x>0\\ -1&x<0\\ [-1,1]&x=0\end{cases}.$
It can be checked that $\nabla h(\bar{x})=\mathrm{sign}(\bar{x})$ for any $\bar{x}\neq 0$. There thus exists small $\varepsilon>0$, for any $|x-\bar{x}|<\varepsilon$, such that
$|x|\leq|\bar{x}|+\langle x-\bar{x},\mathrm{sign}(\bar{x})\rangle+\frac{\sigma}{2}\|x-\bar{x}\|^{2},$
for any $\sigma>0$. Moreover, if $\bar{x}=0$, for small $\varepsilon>0$ and $|x-\bar{x}|<\varepsilon$, we can choose $\bar{w}=\mathrm{sign}(x)\in\partial h(\bar{x})|_{\bar{x}=0}$. One can readily verify that
$|x|\leq\langle x,\mathrm{sign}(x)\rangle+\frac{\sigma}{2}\|x-\bar{x}\|^{2},$
for any $\sigma>0$. For the case $x\in\mathbb{R}^{n}$, the subdifferential of $h(x)$ is the set of all subgradients that can be generated as follows
$\partial h(x)$ =$=\big{\{}g\mid\|g\|_{\infty}\leq 1,\ g^{T}x=\|x\|_{1}\big{\}}.$
For $\bar{x}=\left(\bar{x}_{1},\bar{x}_{2},\ldots,\bar{x}_{n}\right)$, $x\in B_{\varepsilon}(\bar{x})$ with small enough $\varepsilon$, we can choose
$\bar{w}=\left(g_{1},g_{2},\ldots,g_{n}\right),\quad\begin{cases}g_{i}=\nabla|\bar{x}_{i}|=\mathrm{sign}(\bar{x}_{i}),\ &\text{if}\ \bar{x}_{i}\neq 0,\\ g_{i}=\mathrm{sign}(x_{i}),\ &\text{if}\ \bar{x}_{i}=0.\end{cases}$
It can be verified that
$\sum_{i=1}^{n}{\left|x_{i}\right|\leq\sum_{i=1}^{n}{\left|\bar{x}_{i}\right|}+\sum_{i=1}^{n}{\left(x_{i}-\bar{x}_{i}\right)\cdot g_{i}}+\frac{\sigma}{2}\sum_{i=1}^{n}{\left(x_{i}-\bar{x}_{i}\right)^{2}}},$ (4.22)
for any $\sigma>0$. We thus conclude that the anisotropic $l_{1}$-norm satisfies (4.21). ∎
@@
$H\left(x,w,\bar{y},y\right)=f\left(x\right)+\left<x,w\right>-g_{1}^{*}\left(w\right)-g_{2}\left(x\right)+\frac{\sigma}{2}\left\|x-\bar{y}\right\|^{2}+\frac{1}{2}\left\|\bar{y}-y\right\|_{M}^{2},$ (4.23)
for proving the global convergence of the Algorithm 1 for the case $g_{1}$ being non-smooth.
@@
Let $\left\{x^{n}\right\}_{n}$ is generated from $npDCAe_{nls}$ in Algorithm 1 for solving (1.1). Suppose that $\left\{x^{n}\right\}_{n\in N}$ has a cluster $x^{*}$ and $H\left(x,w,\bar{y},y\right)$ is a KL function, $N_{0}=\max(n_{0},n_{1})$, then the following statements hold
@@
    Monotonicity and boundedness: The sequence $\left\{H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)\right\}_{n\geq N_{0}}$ is decreasing monotonically and converges to some limits: $\lim_{n\rightarrow\infty}H\left(x^{n+1},\bar{w}^{n},\bar{x}^{n},x^{n}\right)=\bar{\zeta}$. Moreover, the iteration sequence $\{x^{n}\}_{n}$ is bounded.
@@
(i). Using the conjugate property $\bar{x}^{n},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)=g_{1}\left(\bar{x}^{n}\right)$, we derive
$x^{n+1},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)=\left<x^{n+1}-\bar{x}^{n},\bar{w}^{n}\right>+g_{1}\left(\bar{x}^{n}\right).$ (4.24)
For convenience, we define an auxiliary and intermediate function
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right):=f\left(x^{n+1}\right)+\left<x^{n+1},\bar{w}^{n}\right>-g_{1}^{*}\left(\bar{w}^{n}\right)-g_{2}\left(x^{n+1}\right).$ (4.25)
First with (4.25), then by (4.20) with $\bar{x}=\bar{x}^{n}$ and $x=x^{n+1}$, for any $\bar{w}^{n}\in\partial g_{1}(\bar{x}^{n})$, we have
$F\left(x^{n+1},\bar{w}^{n},x^{n}\right)$ $=f\left(x^{n+1}\right)+\left<x^{n+1}-\bar{x}^{n},\bar{w}^{n}\right>+g_{1}\left(\bar{x}^{n}\right)-g_{2}\left(x^{n+1}\right)$
$\leq f\left(x^{n+1}\right)+g_{1}\left(x^{n+1}\right)-g_{2}\left(x^{n+1}\right)-\frac{\mu}{2}\left\|x^{n+1}-\bar{x}^{n}\right\|^{2},$
where the equality comes from (4.24). Using the energy decreasing (4.1), together with the line search condition (3.9), for the case 1 of (3.6), we obtain

### v2

@@ 4 Convergence analysis
This section analyzes the convergence of the two proposed algorithms and addresses a generalized special case. We first present the convergence for the case that $f(x)$ and $g_{1}(x)$ have Lipschitz gradients while $g_{2}$ is non-smooth. For the case $g_{1}$ being non-smooth, we rigorously prove convergence for the scenario where $g_{1}(x)$ is non-differentiable, and $g_{2}(x)$ is Lipschitz differentiable. These two cases can cover the most practical computational scenarios encountered in applications.
@@ 4.1 Convergence analysis of Algorithm 1
Let $\left\{\bar{x}^{n}\right\}_{n}$ be generated by (3.8), then the following inequality holds
@@
As $\bar{x}^{n}$ is a minimizer of the strongly convex function in (3.8), from the first order optimality condition, we obtain
$0\in-\xi^{n}+M(\bar{x}^{n}-y^{n})+\partial f(\bar{x}^{n})+\partial g_{1}(\bar{x}^{n}).$ (4.2)
By convexity of $f$ and $g_{1}$, for any $x^{n}$, we have
$f(\bar{x}^{n})+\langle s_{f},x^{n}-\bar{x}^{n}\rangle\leq f(x^{n}),\quad\forall s_{f}\in\partial f(\bar{x}^{n}),$
$g_{1}(\bar{x}^{n})+\langle s_{g_{1}},x^{n}-\bar{x}^{n}\rangle\leq g_{1}(x^{n}),\quad\forall s_{g_{1}}\in\partial g_{1}(\bar{x}^{n}),$
which leads to the following inequality by noting $\xi^{n}-M(\bar{x}^{n}-y^{n})\in\partial f(\bar{x}^{n})+\partial g_{1}(\bar{x}^{n})$ from (4.2)
$\langle\xi^{n}-M(\bar{x}^{n}-y^{n}),x^{n}-\bar{x}^{n}\rangle+f(\bar{x}^{n})+g_{1}(\bar{x}^{n})\leq f(x^{n})+g_{1}(x^{n}).$
By the expansion of $\frac{1}{2}\|x^{n}-y^{n}\|_{M}^{2}=\frac{1}{2}\|(x^{n}-\bar{x}^{n})+(\bar{x}^{n}-y^{n})\|_{M}^{2}$, we obtain
$\frac{1}{2}\|\bar{x}^{n}-y^{n}\|_{M}^{2}+\langle M(\bar{x}^{n}-y^{n}),x^{n}-\bar{x}^{n}\rangle+\frac{1}{2}\|\bar{x}^{n}-x^{n}\|_{M}^{2}=\frac{1}{2}\|x^{n}-y^{n}\|_{M}^{2}.$
Adding the above two inequalities gives exactly
$\left<-\xi^{n},\bar{x}^{n}\right>+\frac{1}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+f\left(\bar{x}^{n}\right)+g_{1}\left(\bar{x}^{n}\right)\leq\left<-\xi^{n},x^{n}\right>+\frac{1}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
@@
$+f\left(x^{n}\right)+g_{1}\left(x^{n}\right)-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
@@
Remembering $y^{n}=x^{n}+\beta_{n}(x^{n}-x^{n-1})$, this leads to the inequality (4.1). ∎∎
@@
This leads to (4.4), and the proof is finished. ∎∎
@@
(iii). Given the boundedness of $\left\{x^{n}\right\}_{n}$ from (ii), there exists at least one cluster point $x^{*}$ with a convergent subsequence $\left\{x^{n_{i}}\right\}_{n_{i}}$ satisfying
@@
$M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)-\nabla f\left(\bar{x}^{n_{i}}\right)-\nabla g_{1}\left(\bar{x}^{n_{i}}\right)+\nabla(f+g_{1})({x}^{n_{i}})\in-\partial g_{2}({x}^{n_{i}})+\nabla(f+g_{1})({x}^{n_{i}}).$
@@
$M\left(\bar{x}^{n_{i}}-y^{n_{i}}\right)+(\nabla f({x}^{n_{i}})-\nabla f(\bar{x}^{n_{i}}))+(\nabla g_{1}({x}^{n_{i}})-\nabla g_{1}(\bar{x}^{n_{i}})).$ (4.14)
@@
Setting $\tilde{L}=L+L_{g_{1}}$, according to $\bar{x}^{n}-x^{n}=\frac{\lambda_{n}}{1+\lambda_{n}}\left(x^{n+1}-x^{n}\right)$, $\bar{x}^{n-1}-x^{n-1}=\frac{\lambda_{n-1}}{1+\lambda_{n-1}}\left(x^{n}-x^{n-1}\right)$, we obtain the following formulation
@@
$\geq\psi\prime\left(A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-\zeta\right)\cdot\left[A\left(x^{n},\bar{x}^{n-1},x^{n-1}\right)-A\left(x^{n+1},\bar{x}^{n},x^{n}\right)\right]\cdot$
@@
Finally, we obtain $\sum_{n=1}^{\infty}{\left\|x^{n}-x^{n-1}\right\|}<\infty$. ∎∎
@@ 4.2 Convergence analysis of Algorithm 2
The following Lemma 3, Lemma 4, and Theorem 2 are obtained on this basis, and we will not write them out one by one later on. Following the analytical framework established for Algorithm 2, we initially demonstrate that Lemma 3 serves to rigorously characterize the quasi-descent property intrinsic to the energy functional governing the system.
@@
Let $\{\bar{x}^{n}\}_{n}$ is generated by (3.10), then the following inequality holds
@@
$\left<\nabla f\left(y^{n}\right)-\xi^{n},y\right>+\frac{L}{2}\left\|y-y^{n}\right\|_{M}^{2}+g_{1}\left(y\right).$
@@
$\left<\nabla f\left(y^{n}\right)-\xi^{n},\bar{x}^{n}\right>+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(\bar{x}^{n}\right)\leq\left<\nabla f\left(y^{n}\right)-\xi^{n},x^{n}\right>+\frac{L}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
@@
$\left<\nabla f\left(y^{n}\right),\bar{x}^{n}-x^{n}\right>+\frac{L}{2}\left\|\bar{x}^{n}-y^{n}\right\|_{M}^{2}+g_{1}\left(\bar{x}^{n}\right)-g_{2}\left(\bar{x}^{n}\right)\leq\frac{L}{2}\left\|x^{n}-y^{n}\right\|_{M}^{2}$
@@
Applying the matrix dominance $M\succeq I$ and $y^{n}=x^{n}+\beta_{n}(x^{n}-x^{n-1})$, we derive (4.19).∎∎
@@ 4.3 Convergence analysis for the special case with nonsmooth $g_{1}$
for any $\bar{x}\in\operatorname{dom}g_{1}$ and $\tilde{w}\in\partial g_{1}(\bar{x})$ and $\mu$ can be zero while $g_{1}$ is not strongly convex. We propose a more general case which can cover the widely used case that $g_{1}(x)$ is non-smooth, along with $\nabla g_{2}(x)$ being Lipschitz continuous with constant $L_{g_{2}}$, which includes some important regularization such as SCAD as discussed in [1].
@@
$\widetilde{H}\left(\bar{x},x\right)=E\left(\bar{x}\right)+\frac{1}{2}\left\|\bar{x}-x\right\|_{M}^{2}=f\left(\bar{x}\right)+g\left(\bar{x}\right)+\frac{1}{2}\left\|\bar{x}-x\right\|_{M}^{2},$ (4.22)
for proving the global convergence of the Algorithm 1 for the case $g_{1}$ being non-smooth. For the KL analysis with $\partial\widetilde{H}(\bar{x}^{n},x^{n})$, it will turn out that the most challenging problem is caused by $\partial g_{1}(\bar{x}^{n})$. With $\tilde{H}(\bar{x},x)$, one can replace $\partial g_{1}(\bar{x}^{n})$ through the optimality condition (3.2).
@@
Let $\left\{x^{n}\right\}_{n}$ is generated from $npDCAe_{nls}$ in Algorithm 1 for solving (1.1). Supposing that $\left\{x^{n}\right\}_{n}$ has a cluster $x^{*}$ and $\widetilde{H}\left(\bar{x},x\right)$ is a KL function, $N_{0}=\max\mathrm{(}n_{0},n_{1})$, the following statements hold
@@
    Monotonicity and boundedness: The sequence $\left\{\widetilde{H}\left(\bar{x}^{n},x^{n}\right)\right\}_{n\geq N_{0}}$ is decreasing monotonically and converges to some limits: $\lim_{n\rightarrow\infty}\widetilde{H}\left(\bar{x}^{n},x^{n}\right)=\bar{\zeta}$. Moreover, the iteration sequence $\{x^{n}\}_{n}$ is bounded.
@@
    Optimality: Any cluster point of $\left\{x^{n}\right\}_{n}$ is a critical point of the problem (1.1), i.e. it satisfies the first order optimality condition.
@@
(i). From Lemma 1 we derive
$E\left(\bar{x}^{n}\right)\leq E\left(x^{n}\right)+\frac{\beta_{n}^{2}}{2}\left\|x^{n}-x^{n-1}\right\|_{M}^{2}-\frac{1}{2}\left\|\bar{x}^{n}-x^{n}\right\|_{M}^{2}.$
Due to the non-monotone line search, we have
$E\left(x^{n}\right)\triangleq E\left(\bar{x}^{n-1}+\lambda_{n-1}d^{n-1}\right)\leq E\left(\bar{x}^{n-1}\right)-\eta\lambda_{n-1}\left\|d^{n-1}\right\|^{2}+\nu_{n-1}.$